OpenStudy (anonymous):

solve ( sq rt of x+3) = 20 solve for x

OpenStudy (mertsj):

Square both sides. first.

OpenStudy (mertsj):

\[x+3=400\]

OpenStudy (mertsj):

\[x=397\]

OpenStudy (anonymous):

thats what thought but i wsnt sure

OpenStudy (mertsj):

Good for you.

OpenStudy (anonymous):

how do u do this one then (sq rt of 2x+5) = (sq rt of x + 12)

OpenStudy (mertsj):

Same thing. Square both sides

OpenStudy (mertsj):

\[2x+5=x+12\]

OpenStudy (mertsj):

\[x=7\]

OpenStudy (anonymous):

how did u get 7

OpenStudy (mertsj):

Subtract x from both sides and then subtract 5 from both sides.

OpenStudy (anonymous):

can u draw it out for me

OpenStudy (mertsj):